OpenStudy (anonymous):

Decide whether the given x value is a zero of the function. f(x) = x3 + 5x2 + x + 5; x = –5 Select one: a. No, because there is more than one zero. b. No, because substituting the value does/does not make the function equal to zero. Incorrect c. Yes, because substituting the value does/does not make the function equal to zero. d. none of the above

OpenStudy (anonymous):

@ganeshie8 @thomaster

OpenStudy (ikram002p):

your funtion is \(\large f(x) = x^3 + 5x^2 + x + 5\) right ? x=-5 is a function zero \(\iff \) f(-5) =0 so check \(\large f(-5) = (-5)^3 + 5(-5)^2 + (-5) + 5\) if its equal to zero , then x=-5 is a function zero .

OpenStudy (anonymous):

so which one is it

OpenStudy (ikram002p):

-.- try to check and see , whats the value of f(-5) ?!

OpenStudy (anonymous):

I put it in http://www.wolframalpha.com/input/?i=f%28-5 and it says its even

OpenStudy (anonymous):

oh its 0

OpenStudy (anonymous):

I know the answer isn't b

OpenStudy (ikram002p):

http://www.wolframalpha.com/input/?i=+%28-5%29^3+%2B+5%28-5%29^2+%2B+%28-5%29+%2B+5 ok good then whats the answer ?

OpenStudy (anonymous):

x=-5 is a zero of the function since it satisfies the function when we substitute\[f \left( -5 \right)=0\]. answer will be option c

OpenStudy (anonymous):

ok thanks